# A Systematic Literature Review on Vehicular Collaborative Perception – A Computer Vision Perspective

### v1

@@ Abstract
Index Terms:
Autonomous Driving, Connected Autonomous Vehicles, Cooperative-Intelligent Transportation Systems, Computer Vision, Collaborative Perception, Collective Perception.
@@ I-A Related Work
TABLE I: Summary of Surveys in Vehicular Collaborative Perception. Mod.: Modality, Co.: Collaborative type, OD: Object Detection, OT: Object Tracking, MP: Motion Prediction, SS: Semantic segmentation, LD: Lane detection, MT/TA: Multi-task.Task agnostic, LE: Localization error, TL: Time latency, CB: Communication bandwidth constraint, CI: Communication Interruption, DS: Domain Shift, Hetero.: Heterogeneous system, Adv.: Adversarial attack, DA: Dataset, ES: Evaluation scenarios, EM: Evaluation Metrics, AS: Ablation Study
@@ II-A2 Search and Selection
Figure 5: Procedure of the search and selection
@@ II-B Metadata Analysis
This section presents the metadata analysis conducted to identify research trends in cooperative perception. Figure 4 visualizes the number of publications from 2019 to 2024, showing a steady increase, particularly in 2023. This growth reflects maturing research foundations, expanding real-world applications, and strong funding support. It also indicates growing interest among researchers in the field of CP, suggesting the potential for exponential growth in the future. However, due to delays in updating records, such as those from IEEE Xplore, and the cutoff date for paper collection being March 2024, the actual number of publications for the most recent year is likely higher than reported. Figure 6 presents the distribution of publications by region, highlighting that the primary contributors are from Asia and North America, with 54 and 38 publications. Their dominance is likely driven by rapid advancements in V2X technology, including 5G and connected vehicle infrastructure, supported by strong government policies and funding. Although Europe has fewer publications, it still makes valuable contributions in areas like standardization and foundational research. Additionally, Figure 7 depicts the distribution of publications across journals and conferences, with ICRA, CVPR, and IEEE T-IV as the top venues. This distribution highlights the evolution of CP research beyond its origins in communication and networking, expanding into computer vision, AI, and robotics. The increasing presence in conferences such as CVPR, NeurIPS, and ICCV highlights the growing influence of deep learning and visual perception, while robotics-focused venues like ICRA and IEEE RA-L emphasize CP ’s integration into autonomous systems. This trend demonstrates how advancements in these fields are driving CP forward and provides valuable insights for researchers seeking impactful publication venues.
Figure 6: Number of publications by region. NA represents North America
Figure 7: Number of publications per venue (Top 10)
@@ IV Modality Type (RQ1)
In this section, we provide an in-depth examination of the different modalities in CP. Our systematic review identifies three primary modalities in the reviewed literature: LiDAR, camera, and their combination. Figure 8 illustrates the distribution of CP publications by modality, highlighting that LiDAR-based CP is the predominant focus, represented by 63 studies. In comparison, Camera-based CP and LiDAR-Camera-based CP are less explored, with only 13 and 12 studies. The widespread adoption of LiDAR in CP research can be attributed to its ability to directly capture high-precision 3D spatial information, making it particularly effective for 3D object detection. In contrast, camera-based CP relies on depth estimation, which inherently introduce uncertainties and often result in reduced accuracy for 3D perception tasks. Although LiDAR-camera fusion has the potential to leverage the strengths of both modalities, combining the rich texture from cameras with the accurate depth measuring of LiDAR, its adoption remains limited. This is likely due to the increased complexity of data fusion, which pose significant challenges. Other modalities, such as radar, infrared cameras, and event cameras, are notably absent from peer-reviewed papers that meet the selection criteria, indicating a potential research gap and an opportunity for further exploration.
Figure 8: Number of publications on CP by modality: LiDAR, Camera, and LiDAR-Camera.
@@ V Collaboration type (RQ1)
This section presents an in-depth review of the various collaboration types in CP: Early, Intermediate, Late, and Hybrid. As shown in Figure 10, research in this field is dominated by the Intermediate collaboration type, with 71 studies focused on this approach. In comparison, the Early, Late, and Hybrid types are less represented, with 6, 15, and 6 studies, respectively. Intermediate collaboration is preferred for its balance between communication efficiency and perception accuracy, as it transmits processed features rather than raw sensor data or final results. In contrast, the limited adoption of Early, Late, and Hybrid collaboration reflects challenges in bandwidth constraints, synchronization, information loss, and fusion complexity, emphasizing the need for adaptive collaboration strategies to improve efficiency.
Figure 10: Number of publications on CP by Collaboration type: Early, Intermediate (Inter.), Late, and Hybrid.
@@ V-B2 Attention-Based Feature Fusion
The attention mechanism [82] is effective in capturing long-range dependencies and contextual relationships, making it highly suitable for feature weighting during fusion. For example, Wang et al. propose the F-Transformer [47], a point cloud fusion transformer that employs only Transformer encoder to fuse features from different views. Unlike conventional Transformers, it omits position embeddings since the spatial arrangement of views is arbitrary, and there is no inherent ordering relationship between features from multiple perspectives. This design enhances robustness by preventing the model from making erroneous spatial assumptions and instead allowing it to focus on learning meaningful feature correlations across views. Xu et al. introduce the V2X-ViT [16], designed to fuse information across on-road agents efficiently. It enhances self-attention by incorporating an additional weight matrix tailored to the type of the source and target agents. For example, agents are categorized as either vehicles or infrastructure, and the weight matrix dynamically adjusts to optimize collaboration based on their type. Hu et al. [75] introduce a spatial confidence-aware attentive fusion, where a spatial confidence map identifies perceptual uncertainty across different areas, serving as a basis for attention learning. This method prioritizes features with higher confidence during fusion, enhancing reliability. Lu et al. [55] propose a robust multiscale attentive fusion to mitigate noise from spatial misalignment. This method leverages features at different scales: finer scales provide detailed semantic information, while coarser scales offer robustness against spatial noise, thus maintaining semantic density and enhancing overall robustness. Yang et al. [52] address temporal noise using the spatial-temporal collaboration transformer (STCFormer), which features decoupled spatial and temporal cross-attention. STCFormer follows the architecture of a vanilla transformer but incorporates three customized modules: temporal cross-attention, decoupled spatial attention, and adaptive late fusion. The temporal cross-attention captures historical context across agents to enhance the representation of the current frame, mitigating point cloud sparsity caused by fast-moving objects. The decoupled spatial attention fuses spatial features from multiple agents, while the adaptive late fusion module integrates spatial features using weight maps. With these customized modules, STCFormer achieves robust detection performance even in dynamic environments. Despite its effectiveness, the computational complexity of attention mechanisms $O(N^{2})$ poses scalability challenges. To address this, Yang et al. [56] utilized a deformable cross-attention module that selectively focuses on informative locations, significantly reducing computational demands and memory usage. Unlike standard attention, which assigns weights to all elements in the feature space, deformable attention selectively attends to a sparse set of informative locations, improving computational efficiency and scalability.
@@ V-B3 Graph-Based Feature Fusion
Graph structures are practical tools to represent complex relationships among data elements, where features are modeled as nodes interconnected by edges. These edges depict interactions between features. Graph Neural Network (GNN) [83] is adept at processing this graph-structured data, facilitating effective information aggregation and propagation across the network. Multi-agent collaboration, for instance, can be conceptualized as a graph where nodes represent individual agents and edges represent inter-agent collaborations. Wang et al. [84] introduce a spatially-aware GNN where each agent maintains a local graph with nodes holding state representations. These states are updated via a trainable neural network such as ConvGRU, which processes edge-weighted feature maps from all nodes to output updated node representations. This method incorporates historical context, enhancing temporal alignment and enabling joint object detection and motion prediction. Li et al. [45] propose a collaboration graph with trainable edge weights reflecting the collaboration strength between agents. These spatially and temporally aware weights allow agents to identify regions requiring collaboration dynamically. Xiang et al. [22] further evolve this concept by introducing the H3GAT, a heterogeneous 3D graph attention model that integrates attention mechanisms with GNNs. This model captures local and global interactions, preserving detail and providing a comprehensive context. Liu et al. [58] employ a multiscale graph-attention technique to extract more comprehensive semantic information across different levels of granularity, enhancing feature integration.
@@ VI Perception Tasks (RQ1)
There are various critical perception tasks that can benefit from a collaborative approach, including object detection, object tracking, motion prediction, semantic segmentation, and lane detection. Object Detection (OD) identifies and locates objects within a sensor frame, establishing a foundation for further perception processes. Object Tracking (OT) involves monitoring the dynamic status of an object across multiple frames, while Motion Prediction (MP) aims to forecast the future movements or intentions of an object. Semantic Segmentation (SS) plays a crucial role in scene understanding, helping CAVs identify drivable areas and provide essential information for subsequent tasks. Lane Detection (LD) is integral to determining road boundaries and lane markings, enabling CAVs to comprehend the geometry of the road network. This section provides a comprehensive overview of collaborative methods used in detection, tracking, motion prediction, semantic segmentation, and lane detection. Additionally, it introduces concepts of Multi-Task (MT) and Task-agnostic (TA) pipelines, which are pivotal in enhancing the efficiency and accuracy of vehicle perception systems. Figure 11 presents the distribution of publications by task, indicating a substantial focus on OD, which accounts for 78 studies. In contrast, the number of papers on SS, OT, and MP is comparatively low, with only 6, 5, and 5 studies, respectively. This disparity highlights the dominant research interest in object detection, while other tasks remain underexplored, suggesting potential areas for further investigation.
@@
Figure 11: Number of publications on CP by Task: Object Detection (OD), Semantic Segmentation (SS), Object Tracking (OT), Motion Prediction (MP), Lane Detection (LD), Multi-Task (MT), and Task-agnostic (TA).
@@ VI-A Collaborative Object Detection
Object detection is a fundamental perception task that focuses on identifying and locating relevant objects from raw sensor data. Typically, object detection results are presented as bounding boxes, each labeled with the corresponding object category. These bounding boxes can vary in representation: they may appear in 2D, BEV, or 3D formats. 2D bounding boxes, often used in camera-based 2D object detection, capture object on image plane. BEV representation disregards height and emphasizes the spatial layout of dynamic objects on the road plane, which is often sufficient for downstream tasks such as planning. The 3D format includes height and z-axis position, offering a more comprehensive view of the scene. This section discusses COD across these different representations, with 3D being the most prevalent form in COD applications. All papers on COD that meet our criteria are summarized in Tables XXVIII and XXIX.
@@ VI-B Collaborative Semantic Segmentation
Semantic segmentation is a process designed to assign a semantic class label to every pixel in an image or every point in a LiDAR scan. This technique offers a granular understanding of scenes, going beyond object detection that typically uses bounding boxes to localize objects. Semantic segmentation facilitates the precise delineation of object boundaries and enables the identification of multiple instances within the same scene. However, visual occlusions can create areas where semantic labels cannot be accurately predicted. Through V2X collaboration, CAVs can extend their FOV and supplement the semantic labels of occluded areas, thus achieving a more comprehensive understanding of their surroundings. This section summarizes and categorizes Collaborative Semantic Segmentation (CSS) approaches based on their representation format. All papers on CSS that meet the selection criteria are listed in Table VIII.
TABLE VIII: Overview of the methods for Collaborative Semantic Segmentation (CSS). V: Vehicle, I: Infrastructure, UAV: Unmanned Aerial Vehicle, Raw: Raw data fusion, Trad Feat: Traditional Feature Fusion, Atten Feat: Attention Feature Fusion.
@@ VI-C Collaborative Object Tracking
Object tracking involves locating and following object trajectories across sequences of video frames or point cloud data. Accurate tracking enables determination of an object’s position, velocity, and acceleration, collectively understood as its motion status. Challenges in object tracking, such as dynamic changes in appearance, occlusions, and complex motion patterns, necessitate robust algorithms for continuous and precise tracking. Multi-view collaboration is a promising solution to address occlusions and maintain continuous tracking. This section categorizes collaborative object tracking into two approaches: tracking with Collaborative Object Detection (COD)and tracking without COD. Tracking with COD integrates closely with collaborative detection outcomes, enhancing subsequent perception processes. Alternatively, tracking without COD offers flexibility by fusing perception results from multiple agents independently. Both methods predominantly utilize Kalman filters and their variants for tracking and incorporate uncertainty propagation to refine their tracking processes. All papers on Collaborative Object Tracking (COT) that meet the selection criteria are summarized in Table IX.
TABLE IX: Overview of the methods for collaborative object tracking (COT). V: Vehicle, I: Infrastructure, Obj: Object, COD: Collaborative object detection, Atten Feat: Attention Feature Fusion, Obj Fusion: Object-level Fusion.

### v3

@@ I-A Related Work
TABLE I: Summary of Surveys in Vehicular Collaborative Perception. Mod.: Modality, Co.: Collaborative type, OD: Object Detection, OT: Object Tracking, MP: Motion Prediction, $S$: Semantic segmentation, LD: Lane detection, MT/TA: Multi-task.Task agnostic, LE: Localization error, TL: Time latency, CB: Communication bandwidth constraint, CI: Communication Interruption, DS: Domain Shift, Hetero.: Heterogeneous system, Adv.: Adversarial attack, DA: Dataset, ES: Evaluation scenarios, EM: Evaluation Metrics, AS: Ablation Study
@@ II-A2 Search and Selection
Figure 5: Procedure of the search and selection, starting from 4876 items, reduced to 3980 after duplicates were removed, 249 after screening and snowballing, and resulting in 106 studies included in the final review.
@@ II-B Metadata Analysis
This section presents the metadata analysis conducted to identify research trends in cooperative perception. Figure 4 visualizes the number of publications from 2019 to 2024, showing a steady increase, with a pronounced surge in 2023. This trajectory reflects both the maturation of foundational research and the rapid expansion of real-world applications, supported by significant funding initiatives in intelligent transportation and autonomous driving. It also highlights that cooperative perception (CP) is transitioning from an emerging topic to a consolidated research domain. Nevertheless, due to indexing delays (e.g., IEEE Xplore updates) and the cutoff date for data collection in March 2024, the actual number of recent publications is likely higher than reported, suggesting that this upward trend is even stronger than captured here.
TABLE VIII: Summary of surveyed collaborative perception studies across multiple dimensions.
Dimension Categories (with counts) Key observations
Region Asia (54), North America (38), Europe (13), Africa (1) Research is predominantly conducted in Asia and North America.
Publication venue (top 10) ICRA (16), CVPR (8), IEEE T-IV (8), NeurIPS (8), ICCV (5), IEEE RA-L (5), IEEE ITSC (5), IEEE T-ITS (4), IEEE IoTJ (4), IEEE IV (4) Publications are concentrated in leading robotics and vision venues, with fewer in transportation-focused outlets.
Modality LiDAR (63), Camera (13), LiDAR–Camera (12) LiDAR-based approaches remain predominant, reflecting their reliability for accurate 3D detection.
Collaboration Early (6), Intermediate (71), Late (15), Hybrid (6) Intermediate fusion is the predominant strategy, balancing bandwidth efficiency with information richness.
Task Object detection (78), Semantic segmentation (6), Object tracking (5), Motion prediction (3), Lane detection (2), Multi-task (3), Task-agnostic (8) Object detection dominates the field, while other tasks remain comparatively underexplored.
Regional distribution. Table VIII shows that Asia (54) and North America (38) dominate the research landscape. This concentration reflects the substantial investment in V2X testbeds, 5G infrastructure, and large-scale smart mobility projects in countries such as China and the United States. Europe, while contributing fewer studies (13), plays an important role in standardization and cross-border research initiatives (e.g., ETSI standards for cooperative ITS). By contrast, contributions from other regions, including Africa (1), are minimal, underscoring the geographical imbalance of current CP research and highlighting opportunities for more globally distributed investigations.
Publication venues. The majority of papers are published in premier robotics and computer vision venues, including ICRA (16), and CVPR (8) as summarized in Table VIII. This distribution indicates that CP has expanded beyond its original roots in communication and networking, and is increasingly recognized as a computer vision and AI challenge. The growing presence in CVPR, NeurIPS, and ICCV emphasizes the centrality of deep learning and visual perception methods, while robotics-oriented outlets such as ICRA and IEEE RA-L demonstrate the integration of CP into embodied autonomous systems. The shift of publication venues therefore illustrates both methodological diversification and the convergence of perception, AI, and robotics communities around CP.
Modalities. LiDAR-based approaches account for the majority of studies (63), with only 13 camera-only and 12 LiDAR–Camera fusion papers (Table VIII). LiDAR’s predominance stems from its robustness in capturing precise 3D spatial geometry, which is critical for reliable detection in cluttered and occluded environments. Camera-based methods, by contrast, face inherent challenges in depth estimation and performance degradation under illumination changes. LiDAR–Camera fusion remains underrepresented despite its potential to combine complementary strengths (texture-rich visual data and precise depth). This limited adoption reflects the technical challenges of spatial–temporal calibration and multimodal fusion complexity. Moreover, the near absence of alternative modalities (e.g., radar, thermal, event cameras) underscores a critical research gap and signals promising directions for future multimodal CP.
Collaboration strategies. Table VIII further reveals that intermediate collaboration dominates (71 studies), while early (6), late (15), and hybrid (6) approaches are far less common. The predominance of intermediate fusion reflects its ability to balance bandwidth efficiency with perception accuracy by exchanging processed features rather than raw data or final outputs. The limited adoption of other schemes underscores the persistent technical barriers: early fusion faces prohibitive communication demands, late fusion suffers from information loss, and hybrid designs increase synchronization and integration complexity. These findings suggest that while intermediate collaboration is currently the most practical solution, advancing adaptive and flexible collaboration schemes will be crucial for future large-scale deployments.
Tasks. Finally, research efforts are heavily skewed toward object detection (78 studies), with relatively few works on semantic segmentation (6), object tracking (5), and motion prediction (3), as shown in Table VIII. The focus on object detection underscores its central role as a prerequisite for higher-level reasoning, yet the limited attention to other tasks reveals important gaps. In particular, tracking and prediction are essential for safety-critical decision-making, and their underrepresentation highlights opportunities for future work. Similarly, multi-task and task-agnostic designs could support more integrated perception pipelines, but remain at an early stage of development.
Overall, this metadata analysis demonstrates that while cooperative perception has become a rapidly growing and increasingly multidisciplinary field, current research exhibits clear imbalances across regions, modalities, collaboration schemes, and perception tasks. These insights provide a foundation for identifying gaps and charting future research directions.
@@ IV Modality Type (RQ1)
In this section, we provide an in-depth examination of the different modalities in CP. Our systematic review identifies three primary modalities in the reviewed literature: LiDAR, camera, and their combination.
@@ V Collaboration type (RQ1)
This section presents an in-depth review of the various collaboration types in CP: Early, Intermediate, Late, and Hybrid.
@@ V-B2 Attention-Based Feature Fusion
The attention mechanism [82] is effective in capturing long-range dependencies and contextual relationships, making it highly suitable for feature weighting during fusion. For example, Wang et al. propose the F-Transformer [47], a point cloud fusion transformer that employs only Transformer encoder to fuse features from different views. As illustrated in Fig. LABEL:fig:att_fusion, features from multiple entities are represented as tokens and forwarded to the attention module, where contextual relationships are learned to produce a fused feature representation. Unlike conventional Transformers, it omits position embeddings since the spatial arrangement of views is arbitrary, and there is no inherent ordering relationship between features from multiple perspectives. This design enhances robustness by preventing the model from making erroneous spatial assumptions and instead allowing it to focus on learning meaningful feature correlations across views. Xu et al. introduce the V2X-ViT [16], designed to fuse information across on-road agents efficiently. It enhances self-attention by incorporating an additional weight matrix tailored to the type of the source and target agents. For example, agents are categorized as either vehicles or infrastructure, and the weight matrix dynamically adjusts to optimize collaboration based on their type. Hu et al. [75] introduce a spatial confidence-aware attentive fusion, where a spatial confidence map identifies perceptual uncertainty across different areas, serving as a basis for attention learning. This method prioritizes features with higher confidence during fusion, enhancing reliability. Lu et al. [55] propose a robust multiscale attentive fusion to mitigate noise from spatial misalignment. This method leverages features at different scales: finer scales provide detailed semantic information, while coarser scales offer robustness against spatial noise, thus maintaining semantic density and enhancing overall robustness. Yang et al. [52] address temporal noise using the spatial-temporal collaboration transformer (STCFormer), which features decoupled spatial and temporal cross-attention. STCFormer follows the architecture of a vanilla transformer but incorporates three customized modules: temporal cross-attention, decoupled spatial attention, and adaptive late fusion. The temporal cross-attention captures historical context across agents to enhance the representation of the current frame, mitigating point cloud sparsity caused by fast-moving objects. The decoupled spatial attention fuses spatial features from multiple agents, while the adaptive late fusion module integrates spatial features using weight maps. With these customized modules, STCFormer achieves robust detection performance even in dynamic environments. Despite its effectiveness, the computational complexity of attention mechanisms $O(N^{2})$ poses scalability challenges. To address this, Yang et al. [56] utilized a deformable cross-attention module that selectively focuses on informative locations, significantly reducing computational demands and memory usage. Unlike standard attention, which assigns weights to all elements in the feature space, deformable attention selectively attends to a sparse set of informative locations, improving computational efficiency and scalability.
@@
((a))
((b))
Figure 7: Illustration of intermediate feature fusion: (a) Attention-based feature fusion, where features from ego-vehicle, cooperative vehicle, and infrastructure are transformed into tokens and aggregated using attention mechanisms; (b) Graph-based feature fusion, where features are represented as nodes and fused through message passing and node state updates in a GNN.
@@ V-B3 Graph-Based Feature Fusion
Graph structures are practical tools to represent complex relationships among data elements, where features are modeled as nodes interconnected by edges. These edges depict interactions between features. As shown in Fig. LABEL:fig:graph_fusion, multi-agent collaboration can be conceptualized as a graph where nodes represent individual agents and edges represent inter-agent collaborations. Graph Neural Network (GNN) [83] are well suited for processing such graph-structured data, enabling effective message passing and node state updates that facilitate information aggregation and propagation across the network. For instance, Wang et al. [84] introduce a spatially-aware GNN where each agent maintains a local graph with nodes holding state representations. These states are updated via a trainable neural network such as ConvGRU, which processes edge-weighted feature maps from all nodes to output updated node representations. This method incorporates historical context, enhancing temporal alignment and enabling joint object detection and motion prediction. Li et al. [45] propose a collaboration graph with trainable edge weights reflecting the collaboration strength between agents. These spatially and temporally aware weights allow agents to identify regions requiring collaboration dynamically. Xiang et al. [22] further evolve this concept by introducing the H3GAT, a heterogeneous 3D graph attention model that integrates attention mechanisms with GNNs. This model captures local and global interactions, preserving detail and providing a comprehensive context. Liu et al. [58] employ a multiscale graph-attention technique to extract more comprehensive semantic information across different levels of granularity, enhancing feature integration.
@@ VI Perception Tasks (RQ1)
There are various critical perception tasks that can benefit from a collaborative approach, including object detection, object tracking, motion prediction, semantic segmentation, and lane detection. Object Detection (OD) identifies and locates objects within a sensor frame, establishing a foundation for further perception processes. Object Tracking (OT) involves monitoring the dynamic status of an object across multiple frames, while Motion Prediction (MP) aims to forecast the future movements or intentions of an object. Semantic Segmentation (SS) plays a crucial role in scene understanding, helping CAVs identify drivable areas and provide essential information for subsequent tasks. Lane Detection (LD) is integral to determining road boundaries and lane markings, enabling CAVs to comprehend the geometry of the road network. This section provides a comprehensive overview of collaborative methods used in detection, tracking, motion prediction, semantic segmentation, and lane detection. Additionally, it introduces concepts of Multi-Task (MT) and Task-agnostic (TA) pipelines, which are pivotal in enhancing the efficiency and accuracy of vehicle perception systems.
@@ VI-A Collaborative Object Detection
Object detection is a fundamental perception task that focuses on identifying and locating relevant objects from raw sensor data. Typically, object detection results are presented as bounding boxes, each labeled with the corresponding object category. These bounding boxes can vary in representation: they may appear in 2D, BEV, or 3D formats. 2D bounding boxes, often used in camera-based 2D object detection, capture object on image plane. BEV representation disregards height and emphasizes the spatial layout of dynamic objects on the road plane, which is often sufficient for downstream tasks such as planning. The 3D format includes height and z-axis position, offering a more comprehensive view of the scene. This section discusses COD across these different representations, with 3D being the most prevalent form in COD applications. All papers on COD that meet our criteria are summarized in Tables XXIX and XXX.
@@ VI-B Collaborative Semantic Segmentation
Semantic segmentation is a process designed to assign a semantic class label to every pixel in an image or every point in a LiDAR scan. This technique offers a granular understanding of scenes, going beyond object detection that typically uses bounding boxes to localize objects. Semantic segmentation facilitates the precise delineation of object boundaries and enables the identification of multiple instances within the same scene. However, visual occlusions can create areas where semantic labels cannot be accurately predicted. Through V2X collaboration, CAVs can extend their FOV and supplement the semantic labels of occluded areas, thus achieving a more comprehensive understanding of their surroundings. This section summarizes and categorizes Collaborative Semantic Segmentation (CSS) approaches based on their representation format. All papers on CSS that meet the selection criteria are listed in Table IX.
TABLE IX: Overview of the methods for Collaborative Semantic Segmentation (CSS). V: Vehicle, I: Infrastructure, UAV: Unmanned Aerial Vehicle, Raw: Raw data fusion, Trad Feat: Traditional Feature Fusion, Atten Feat: Attention Feature Fusion.
@@ VI-C Collaborative Object Tracking
Object tracking involves locating and following object trajectories across sequences of video frames or point cloud data. Accurate tracking enables determination of an object’s position, velocity, and acceleration, collectively understood as its motion status. Challenges in object tracking, such as dynamic changes in appearance, occlusions, and complex motion patterns, necessitate robust algorithms for continuous and precise tracking. Multi-view collaboration is a promising solution to address occlusions and maintain continuous tracking. This section categorizes collaborative object tracking into two approaches: tracking with Collaborative Object Detection (COD) and tracking without COD. Tracking with COD integrates closely with collaborative detection outcomes, enhancing subsequent perception processes. Alternatively, tracking without COD offers flexibility by fusing perception results from multiple agents independently. Both methods predominantly utilize Kalman filters and their variants for tracking and incorporate uncertainty propagation to refine their tracking processes. All papers on Collaborative Object Tracking (COT) that meet the selection criteria are summarized in Table X.
TABLE X: Overview of the methods for collaborative object tracking (COT). V: Vehicle, I: Infrastructure, Obj: Object, COD: Collaborative object detection, Atten Feat: Attention Feature Fusion, Obj Fusion: Object-level Fusion.